# Towards sub-milliarcsecond astrometric precision using seeing-limited imaging

### v1

@@ 4 Astrometric Catalogue Extraction
The pipelines begin with images that have been bias-subtracted and flat-field corrected. As part of this work, we are using tools from the MATLAB Astronomy & Astrophysics Package [^1] https://github.com/EranOfek/AstroPack(Ofek 2014; Soumagnac & Ofek 2018; Ofek 2019; Ofek et al. 2023b). The first step is to estimate the image background [^2] Using imProc.background.background. Instead of performing source detection in each image, we rely on the reference catalogue employed by KMTNet, which is constructed from a combination of the OGLE-III star catalogue (Szymański et al., 2011) and the DECam Plane Survey catalogue (Schlafly et al., 2018). This composite catalogue reaches a limiting magnitude of approximately 22.
The next step is to perform PSF photometry. The PSF is constructed from bright and isolated stars in the field: For each source, its position is measured using the weighted first moment, and it is shifted to the pixel origin using Whittaker-Shannon interpolation. To avoid introducing artifacts caused by Whittaker-Shannon interpolation on a finite-size stamp, and artifacts caused by nearby stars, we employ a hybrid model to describe the PSF. The hybrid model is divided into two regimes, determined by the radius $\bar{r}$ from the center of the PSF. For the inner part ($r<\bar{r}$), the PSF is represented by the empirical PSF, while for the outer part ($r\mathord{\geq}\bar{r}$), we use a fit to an analytical model. We adopt the Multivariate $t$-distribution (see §A) as the analytic PSF model because it can account for PSF asymmetries.
@@ 5 Astrometric model
where $X(t_{j}),Y(t_{j}),Z(t_{j})$ are Topocentric Barycentric Cartesian coordinates at time $t_{j}$. Notice that for the KMTNet orientation $x\parallel-\alpha$ and $y\parallel\delta$.
@@ 5.2 Affine Transformation
In this work, we restrict the transformation to an affine model, as we operate on small cutouts with a relatively small field of view in which the distortions are expected to be minimal. However, the method described here is general and can accommodate higher-order transformations if needed for larger or more distorted fields. However, this should be done with care, as this may introduce new degeneracies into the model.
@@ 5.3 Differential Chromatic Refraction
Because the DCR-induced shift is typically small (on the order of $\lesssim 10$ mas), it is not practical to fit this effect separately for every source and epoch. Instead, we group sources into color bins and model the DCR within each bin, using a weighted sum to calculate the effective residuals and weight, as described below. Specifically, sources are binned according to their $V-I$ color, using a bin width of 0.5 mag, resulting in six color bins. The impact and correction of this effect are discussed in §8.1 and illustrated in Figures 2 and 3. In §7, we explain in detail the matrix form that we use to fit the DCR effect.
@@ 6 Minimisation
We follow the Astrometric Global Iterative Solution (AGIS, Lindegren et al. 2012), with some modifications, for the astrometry solution and minimisation. Because we have modified the algorithm to include the colour terms and additional detrending steps, we write the algorithm explicitly for our case here.
@@
$\begin{bmatrix}N_{ss}&N_{se}&N_{sc}\\ N_{es}&N_{ee}&N_{ec}\\ N_{cs}&N_{ce}&N_{cc}\end{bmatrix}\begin{bmatrix}u_{s}\\ u_{e}\\ u_{c}\\ \end{bmatrix}=\begin{bmatrix}b_{s}\\ b_{e}\\ b_{c}\end{bmatrix},$ (13)
where the subscripts $s$, $e$, and $c$ stand for derivatives with respect to the source, epoch, and chromatic parameters, respectively. These sub-components are described in §7. One important result is that in the case of a linear model, only $b$ is a function of the free parameters and the residuals. Therefore, we only need to update $b_{s}$, $b_{e}$, and $b_{c}$, keep $u$, and recalculate the residuals in each iteration. The linear systems were solved using the bi-conjugate gradients method [^5] Using the built-in bicg function in MATLAB.. We define a step as a full update of the model parameters (source, epoch, chromatic terms), and an iteration as the internal loop solving the normal equations (Equation 11).
@@ 6.1 Uncertainties and weights
To identify and suppress outliers, we use a similar approach. Specifically, we apply a running median and a median absolute deviation (MAD) filter to the 2D astrometric scatter as a function of magnitude. Sources whose RMS exceeds three times the local MAD plus the local median are flagged as outliers, and their weights are reduced by a factor of ten.
Figure 2: Astrometric residuals as a function of the DCR-leading terms in the x-axis. Left panels show the residuals without DCR correction, while right panels show the residuals after applying the correction. Error bars represent the mean and standard error of the mean (i.e., standard deviation divided by the square root of the number of measurements) in each bin of $V-I$ color and corresponding DCR term. Only the top 33% of sources, based on 2D single-epoch precision ($\leq 10$ mas), are included.
@@
7.1 Source and Epoch Parameters
@@
The epoch model includes an affine transformation that maps reference-frame coordinates to detector coordinates. For a single epoch, the design matrices are:
$H_{e,x}$ $=[x_{0},y_{0},1,0,0,0],$ (17)
$H_{e,y}$ $=[0,0,0,x_{0},y_{0},1],$ (18)
where $x_{0},y_{0}$ are the source coordinates at the reference epoch. The epoch design matrices $H_{e}$ have dimensions $N_{\mathrm{src}}\times 6$, which correspond to the six parameters of the affine transformation on the two axes, as described in Equation 4.
@@
$N_{ee}$ $=H^{\intercal}_{e,x}W_{s}H_{e,x}+H^{\intercal}_{e,y}W_{s}H_{e,y},$ (21)
$b_{e}$ $=H^{\intercal}_{e,x}W_{s}R_{x}+H^{\intercal}_{e,y}W_{s}R_{y},$ (22)
@@
To account for chromatic shifts due to the Earth’s atmosphere, we include a DCR component in the model, as described in 5.3.
@@
yields the DCR coefficients $x_{c}$, which are then used to compute the chromatic correction for all sources within the bin. This model neglects chromatic aberrations due to the telescope optics. The rationale is that, since in most cases, the stars are placed in roughly the same position in the focal plane, then to the first order, such aberrations are roughly constant.
Figure 4: Astrometric residuals as a function of the annual phase for both axes, before (left) and after (right) detrending. Residuals are binned in intervals of 0.05 in phase (fraction of year), and colored by $V-I$ color bins as indicated in the legend. Each point represents the median residual within a bin, and the error bars show the standard error. We use the top 33% best sources, with 2D single-epoch precision of $\leq 10\,\mathrm{mas}$.
Figure 5: Colour–magnitude diagram (CMD) based on Gaia DR3 parameters for a 300$\times$300 pixel cutout of the KMTNet BLG17K0103 field. Colours indicate Gaia DR3 parallax values, with hollow triangles denoting sources with negative parallax.
@@ 7.2 Simple Example
As shown by Lindegren et al. (2012), it is sufficient to neglect the off-diagonal matrices (e.g., $N_{se}$) in each iteration, as long as the parameters are updated sequentially. Based on this argument, we drop the cross-coupling terms and update the source and epoch parameters independently in each step. This allows us to solve the normal equations block by block: first solving for the source parameters using the current epoch solution, and then updating the epoch parameters using the updated source solution.
@@ 8 Detrending
Real data may contain systematic effects that are not included in our initial model. We need a way to estimate these effects and detrend them. To address this, we employ several basic detrending techniques. Specifically, we identify correlations between the astrometric residuals and various parameters, as well as applying the SYSREM algorithm (Tamuz et al. 2005), which operates without requiring a model. Table 2 outlines the workflow for each algorithm and details the application of the detrending procedures. In the following, we present the detrending methods we are using in the order we apply them.
@@ 8.1 Chromatic effects
To visualize the impact of this effect, we run our pipeline without incorporating the parallax and DCR models. The left panels of Figures 2 and 3 illustrate the residuals along the X and Y axes, respectively, binned by color ($V-I$), as a function of the linear terms in the chromatic model (i.e., $AM\sin{q}$ and $AM\cos{q}$), where only sources with a 2D single epoch precision (RMS) of less than 14 mas are presented. Next, the right panels of Figures 2 and 3 show the same after applying the chromatic model. These plots demonstrate that for some stars the chromatic effects may be as large as $\sim 10$ mas, and that the inclusion of this effect in our astrometric model reduces the amplitude of the residuals to below $\approx 1$ mas.
@@ 8.3 Intra-Pixel variation
Intra-pixel variation refers to systematic astrometric shifts that depend on a source’s precise location within a detector pixel. These shifts may arise from various reasons including: sub-pixel sensitivity variations; optical or electronic distortions across the pixel surface (Esplin & Luhman, 2015); or Nyqusit under-sampled PSF (e.g., Ofek 2019).
@@ 8.4 SysRem
SysRem (Tamuz et al., 2005) is a blind iterative detrending algorithm designed to identify and remove systematic effects, originally developed for photometric data. This method works by approximating the residual matrix using the multiplication of two vectors (e.g., using Singular Value Decomposition, or iterative methods). SysRem allows for removing systematic effects whose nature is unknown (a-priori). For example, in photometric data, typically one SysRem vector will act as the zero point per image, while the second vector is a zero point per star (e.g., due to color effects). We never employ SysRem more than once. Otherwise, one can remove real variability from the data.
@@ 9 Step-by-Step Algorithm
As outlined in §6, the pipeline solves a linear model iteratively: we start with an initial solution and then refine the model parameters step by step through repeated cycles. This iterative structure allows us to decouple the various components of the astrometric model. In each iteration, we solve the Normal Equation (see Equation 11) for a specific model component (e.g., proper motion, affine transformation, etc.), update the corresponding parameters, and then proceed to the next component in a fixed order. This modular approach follows the principles of the Astrometric Global Iterative Solution (AGIS) method developed for Gaia astrometry (Lindegren et al., 2012), and facilitates improved convergence as well as reduced parameter correlations. We refer to such an iterative solution of Equation 11 as a ”step”. The first step is always performed without weights. Following this initial step, we perform additional steps to detrend the data and perform another step with weights.
@@
  • SYSREM detrending: After the modelling steps, we optionally run one iteration of the SYSREM algorithm (Tamuz et al., 2005) on the residuals. SYSREM models and removes common-mode systematics across sources and epochs.
@@
[…]
Figure 8: Power spectral density (PSD) of the binned position residuals (corresponding to the orange error bars in Figure 7), shown for the x-axis (top) and y-axis (bottom). The PSD is normalised by the variance of each time series, making the power dimensionless and enabling comparison between sources. The blue line indicates the mean PSD over all sources with 2D single epoch precision of $<10$, while the orange line corresponds to the PSD of the example source shown in Figure 7.
@@ 10 Test on real data
Because our algorithm may be applied in different ways, here we have used algorithm 6 of Table 2. We note that the different algorithmic variations do not result in considerable changes. Figure 6 shows the root-mean-square (rms) of the observed positions of each source around the best-fit astrometric model as a function of $I$-band magnitude. For sources near the scintillation floor ($13.5\lesssim I\lesssim 16$), the typical rms per epoch is $\mathord{\sim}3$ mas in one axis and $\mathord{\sim}5$ mas in both axes. Most of the outliers in Figure 6 (marked with ’asterisks’) are sources with a nearby companion that do not appear in the KMTNet reference catalogue, or are mistreated as a single PSF. Figure 7 provides an example of the residuals as a function of time for a single source in the KMTNet BLG17K0103 field. Figure 8 shows the corresponding power spectral density (PSD) of the residuals, both for this individual star and for the mean over sources with a 2-D single-epoch precision $<10$ mas. In the PSD, there are some indications of red noise presence in the lowest frequencies. This may suggest that there is room for improvement in the current pipeline model and fitting procedure.

### v2

@@ 4 Astrometric Catalogue Extraction
The pipelines begin with images that have been bias-subtracted and flat-field corrected. As part of this work, we are using tools from the MATLAB Astronomy & Astrophysics Package [^1] https://github.com/EranOfek/AstroPack(Ofek2014_MAAT; Soumagnac+Ofek2018_catsHTM; Ofek2019_Astrometry_Code; Ofek+2023PASP_LAST_PipeplineI). The first step is to estimate the image background [^2] Using imProc.background.background. Instead of performing source detection in each image, we rely on the reference catalogue employed by KMTNet, which is constructed from a combination of the OGLE-III star catalogue (szymanski2011optical_OGLE_III_catalog) and the DECam Plane Survey catalogue (schlafly2018decam_catalog). This composite catalogue reaches a limiting magnitude of approximately 22.
The next step is to perform PSF photometry. The PSF is constructed from bright, isolated stars in the field, typically using several tens of stars to ensure a well-sampled empirical PSF. For each source, its position is measured using the weighted first moment, and it is shifted to the pixel origin using Whittaker-Shannon interpolation. Then, we normalize the cutout of each source and set the PSF as the median value in each pixel. To avoid introducing artifacts caused by Whittaker-Shannon interpolation on a finite-size stamp, and artifacts caused by nearby stars, we employ a hybrid model to describe the PSF. The hybrid model is divided into two regimes, determined by the radius $\bar{r}$ from the center of the PSF. For the inner part ($r<\bar{r}$), the PSF is represented by the empirical PSF, while for the outer part ($r\mathord{\geq}\bar{r}$), we use a fit to an analytical model. The optimal transition radius, $\bar{r}$, varies with observing conditions and is typically close to the FWHM. We adopt the Multivariate $t$-distribution (see §A) as the analytic PSF model because it can account for PSF asymmetries.
@@
We adopt a single, spatially invariant PSF per field because each cutout spans $\mathord{\sim}2^{\prime}$ on a side, only $\mathord{\sim}10^{-4}$ of KMTNet’s $4\,\mathrm{deg}^{2}$ frame, over which the spatial PSF gradients (set mainly by smoothly varying optical aberrations) are small. This assumption breaks down for cutouts that span a significant fraction of the telescope’s field of view, where a spatially varying PSF model becomes necessary.
@@ 5 Astrometric model
All astrometric quantities in this work are defined relative to a fiducial reference frame constructed from the ensemble of stars in each field. This formulation follows the fact that our measurements, and those relevant for astrometric microlensing, are differential by construction, i.e., they describe relative displacements of sources with respect to the local reference frame rather than their absolute positions on the sky.
@@
where $X(t_{j}),Y(t_{j}),Z(t_{j})$ are the Barycentric coordinates of Topocentric observer at time $t_{j}$. Notice that for the KMTNet orientation $x\parallel-\alpha$ and $y\parallel\delta$.
In practice, however, we do not fit parallax to the real data analysed in this work, as the annual effect introduces systematics that dominate over the expected parallax signal (see §8.2).
@@ 5.2 Affine Transformation
In this work, we restrict the transformation to an affine model, as we operate on small cutouts with a relatively small field of view in which the distortions are expected to be minimal. However, the method described here is general and can accommodate higher-order transformations if needed for larger or more distorted fields, especially for wider fields of view or higher zenith angle, where atmospheric differential refraction is significant. For our fields of view, this effect is $\mathord{\sim}10^{-2}\,\mathrm{mas}$ (using the formula from seidelmann1992explanatory), so an affine transformation is sufficient. However, this should be done with care, as this may introduce new degeneracies into the model.
@@ 5.3 Differential Chromatic Refraction
Because the DCR-induced shift is typically small (on the order of $\lesssim 10$ mas), it is not practical to fit this effect separately for every source and epoch. Instead, we group sources into color bins and model the DCR within each bin, using a weighted sum to calculate the effective residuals and weight, as described below. Specifically, sources are binned according to their VIV-I color, using a bin width of 0.5 mag, resulting in six color bins. The impact and correction of this effect are discussed in §8.1 and illustrated in Figures 2 and 3. In §7, we explain in detail the matrix form that we use to fit the DCR effect.
@@ 6 Minimisation
We follow the Astrometric Global Iterative Solution (AGIS, lindegren2012astrometric), with some modifications, for the astrometry solution and minimisation. The procedure is mathematically equivalent to the Gauss–Newton method for non-linear least-squares, implemented in block form to separately update source, affine, and DCR parameters. Because we have modified the algorithm to include the colour terms and additional detrending steps, we write the algorithm explicitly for our case here.
@@
$\begin{bmatrix}N_{ss}&N_{se}\\ N_{es}&N_{ee}\\ \end{bmatrix}\begin{bmatrix}u_{s}\\ u_{e}\\ \end{bmatrix}=\begin{bmatrix}b_{s}\\ b_{e}\end{bmatrix}$, (13)
where the subscripts $s$ and $e$ denote derivatives with respect to the source and epoch parameters, respectively. For clarity, we further decompose the per-epoch vector as $\mathbf{e}=(\mathbf{a},\,c)$ when discussing the affine and DCR sub-components in §7. One important result is that in the case of a linear model, only $b$ is a function of the free parameters and the residuals. Therefore, we only need to update $b_{s}$ and $b_{e}$, keep $u$, and recalculate the residuals in each iteration. The linear systems were solved using the bi-conjugate gradients method [^5] Using the built-in bicg function in MATLAB.. We define a step as a full update of the model parameters (source, epoch, chromatic terms), and an iteration as the internal loop solving the normal equations (Equation 11).
@@ 6.1 Uncertainties and weights
To identify and suppress outliers, we use a similar approach. Specifically, we apply a running median and a median absolute deviation (MAD) filter to the 2D astrometric scatter as a function of magnitude. Sources whose RMS exceeds three times the local MAD plus the local median are flagged as outliers, and their weights are reduced by a factor of ten. We adopt this mild down-weighting as a heuristic compromise: it suppresses their influence while keeping them in the fit, because in an iterative scheme, some measurements can switch between inlier and outlier status across iterations.
Figure 2: Astrometric residuals as a function of the DCR-leading terms in the x-axis. Left panels show the residuals without DCR correction, while right panels show the residuals after applying the correction. Error bars represent the mean and standard error of the mean (i.e., standard deviation divided by the square root of the number of measurements) in each bin of VIV-I color and corresponding DCR term. Only the top 33% of sources, based on 2D single-epoch precision ($\leq 10$ mas), are included.
@@
7.1 Source and Affine Parameters
@@
The affine model describes the per-epoch linear transformation that maps reference-frame coordinates to detector coordinates. For a single epoch, the design matrices are:
$H_{a,x}$ $=[x_{0},y_{0},1,0,0,0],$ (17)
$H_{a,y}$ $=[0,0,0,x_{0},y_{0},1],$ (18)
where $x_{0},y_{0}$ are the source coordinates at the reference epoch. The affine design matrices $H_{e}$ have dimensions $N_{\mathrm{src}}\times 6$, which correspond to the six parameters of the affine transformation on the two axes, as described in Equation 4.
@@
$N_{aa}$ $=H^{\intercal}_{a,x}W_{s}H_{a,x}+H^{\intercal}_{a,y}W_{s}H_{a,y},$ (21)
$b_{a}$ $=H^{\intercal}_{a,x}W_{s}R_{x}+H^{\intercal}_{a,y}W_{s}R_{y},$ (22)
@@
To account for chromatic shifts due to the Earth’s atmosphere, we add the per-epoch chromatic term $c$ to the model, as described in §5.3.
@@
yields the DCR coefficients $x_{c}$, which are then used to compute the chromatic correction for all sources within the bin. This model neglects chromatic aberrations due to the telescope optics. The rationale is that, since in most cases, the stars are placed in roughly the same position in the focal plane, then, to the first order, such aberrations are roughly constant.
Figure 4: Astrometric residuals as a function of the annual phase for both axes, before (left) and after (right) detrending. Residuals are binned in intervals of 0.05 in phase (fraction of year), and colored by VIV-I color bins as indicated in the legend. Each point represents the median residual within a bin, and the error bars show the standard error. We use the top 33% best sources, with 2D single-epoch precision of $\leq 10\,\mathrm{mas}$.
Figure 5: Colour–magnitude diagram (CMD) based on Gaia DR3 parameters for a $\times300$ pixel cutout of the KMTNet BLG17K0103 field. Colours indicate Gaia DR3 parallax values, with hollow triangles denoting sources with negative parallax.
@@ 7.2 Simple Example
In this simplified example, the per-epoch parameters reduce to the affine coefficients only, so we denote them by $\mathbf{a}$ and write the normal equations using aa–subscripts.
@@
As shown by lindegren2012astrometric, it is sufficient to neglect the off-diagonal matrices (e.g., $N_{se}$) in each iteration, as long as the parameters are updated sequentially. Based on this argument, we drop the cross-coupling terms and update the source and epoch parameters independently in each step. This allows us to solve the normal equations block by block: first solving for the source parameters using the current epoch solution, and then updating the epoch parameters using the updated source solution.
@@ 8 Detrending
Real data may contain systematic effects that are not included in our initial model. We need a way to estimate these effects and detrend them. To address this, we employ several basic detrending techniques. Specifically, we identify correlations between the astrometric residuals and various parameters, as well as applying the SYSREM algorithm (sysrem_tamuz2005correcting), which operates without requiring a model. Table 2 outlines the workflow for each algorithm and details the application of the detrending procedures. In the following, we present the detrending methods we are using in the order we apply them.
@@ 8.1 Chromatic effects
To visualize the impact of this effect, we run our pipeline without incorporating the parallax and DCR models. The left panels of Figures 2 and 3 illustrate the residuals along the X and Y axes, respectively, binned by color (VIV-I), as a function of the linear terms in the chromatic model (i.e., $AM\sin{q}$ and $AM\cos{q}$), where only sources with a 2D single epoch precision (RMS) of less than 14 mas are presented. Next, the right panels of Figures 2 and 3 show the same after applying the chromatic model. These plots demonstrate that for some stars the chromatic effects may be as large as $\sim 10$ mas, and that the inclusion of this effect in our astrometric model reduces the amplitude of the residuals to below $\approx 1$ mas.
@@ 8.3 Intra-Pixel variation
Intra-pixel variation refers to systematic astrometric shifts that depend on a source’s precise location within a detector pixel. These shifts may arise from various reasons including: sub-pixel sensitivity variations; optical or electronic distortions across the pixel surface (irac_astrometry_esplin2015measuring); or Nyqusit under-sampled PSF (e.g., Ofek2019_Astrometry_Code).
@@ 8.4 SysRem
SysRem (sysrem_tamuz2005correcting) is a blind iterative detrending algorithm designed to identify and remove systematic effects, originally developed for photometric data. This method works by approximating the residual matrix using the multiplication of two vectors (e.g., using Singular Value Decomposition, or iterative methods). SysRem allows for removing systematic effects whose nature is unknown (a-priori). For example, in photometric data, typically one SysRem vector will act as the zero point per image, while the second vector is a zero point per star (e.g., due to color effects). We never employ SysRem more than once. Otherwise, one can remove real variability from the data.
@@ 9 Step-by-Step Algorithm
As outlined in §6, the pipeline solves a linear model iteratively: we start with an initial solution and then refine the model parameters step by step through repeated cycles. This iterative structure allows us to decouple the various components of the astrometric model. In each iteration, we solve the Normal Equation (see Equation 11) for a specific model component (e.g., proper motion, affine transformation, etc.), update the corresponding parameters, and then proceed to the next component in a fixed order. This modular approach follows the principles of the Astrometric Global Iterative Solution (AGIS) method developed for Gaia astrometry (lindegren2012astrometric), and facilitates improved convergence as well as reduced parameter correlations. We refer to such an iterative solution of Equation 11 as a ”step”. The first step is always performed without weights. Following this initial step, we perform additional steps to detrend the data and perform another step with weights.
@@
  • SYSREM detrending: After the modelling steps, we optionally run one iteration of the SYSREM algorithm (sysrem_tamuz2005correcting) on the residuals. SYSREM models and removes common-mode systematics across sources and epochs.
@@
Figure 7: Astrometric residual precision versus source magnitude for different time baselines. For each source, we compute the weighted rms of the residuals about the best-fit astrometric model after binning the residual time series into cadences of 1, 5, 10, and 20 days (colours as indicated). Within each cadence, residuals are averaged per time bin, and the rms of these binned means is calculated with weights proportional to the number of epochs contributing to each bin; only bins containing at least two epochs are retained, and bins with strong outliers are excluded. Thin segments connect measurements of the same source across cadences. Solid lines indicate running median trends in magnitude bins. This short-baseline rms provides an empirical measure of astrometric precision on week–month timescales relevant for astrometric microlensing. Star-level outliers are marked with asterisks.
Figure 8: An example of the astrometric residuals for a source from KMTNet BLG17K0103. The top panel shows the I-band photometric light curve. The middle and bottom panels display the astrometric residuals in the X and Y directions, respectively. In both astrometric panels, the blue dots are the residuals per epoch, while the orange error bars represent the mean residuals in ten-day bins and their standard error of the mean. The 2D astrometric RMS of the epochal measurements is $\mathord{\sim}$5 mas, while the RMS of the orange binned points is 1 mas in both axes.
Figure 9: Power spectral density (PSD) of the binned position residuals (corresponding to the orange error bars in Figure 8), shown for the x-axis (top) and y-axis (bottom). The PSD is normalised by the variance of each time series, making the power dimensionless and enabling comparison between sources. The blue line indicates the mean PSD over all sources with 2D single epoch precision of <10<10, while the orange line corresponds to the PSD of the example source shown in Figure 8.
@@ 10 Test on real data
Because our algorithm may be applied in different ways, here we have used algorithm 6 of Table 2. We note that the different algorithmic variations do not result in considerable changes. Figure 6 shows the root-mean-square (rms) of the observed positions of each source around the best-fit astrometric model as a function of $I$-band magnitude. For sources near the scintillation floor ($13.5\lesssim I\lesssim 16$), the typical rms per epoch is $\mathord{\sim}3$ mas in one axis and $\mathord{\sim}5$ mas in both axes. Most of the outliers in Figure 6 (marked with ’asterisks’) are sources with a nearby companion that do not appear in the KMTNet reference catalogue, or are mistreated as a single PSF. After binning the residual time series in 5–20 day cadences, the bright subset reaches a weighted rms of order $\sim$2 mas per coordinate. This behaviour is shown in Figure 7 and approaches the precision required for astrometric microlensing.
Figure 8 provides an example of the residuals as a function of time for a single source in the KMTNet BLG17K0103 field. Figure 9 shows the corresponding power spectral density (PSD) of the residuals, both for this individual star and for the mean over sources with a 2-D single-epoch precision <10<10 mas. In the PSD, there are some indications of red noise presence in the lowest frequencies. This may suggest that there is room for improvement in the current pipeline model and fitting procedure.